OpenStudy (anonymous):

What are three laws of exponents (for multiplication, division, and powers)? explain each law. Why do we use exponents in scientific notation (you can provide an example)? Write 2 problems containing exponents

OpenStudy (mathstudent55):

1. \(\large a^m \times a^n = a^{m + n} \) 2. \(\large\dfrac{a^m}{a^n} = a^{m - n} \) 3. \(\large (a^m)^n = a^{mn} \)